1. Assume an economy where spending for each sector is:

Household: C = 800 + 0.95Qd

Business: I = 3000

Public: G = 4000, Tr = 7000, Tx = 1000 + 0.3Q

Foreign: X = 1700, Im = 200 + 0.165Q

Solve for:
a) Autonomous Spending
b) Spending Multiplier
c) Disposable Income
d) Consumption Expenditure
e) Household Savings,
f) Imports
g) Net Exports
h) Government Expenditure
i) Budget Deficit

2.) Continuing with the previous problem, what happens to the values in parts a-i if we increase Government Purchases by 1500?

3.) Continuing with 2, what happens if Investment Spending decreased by 500?

4.) Continuing with 3, what happens if exports also decrease by 500?
1
Expert's answer
2020-04-15T08:40:10-0400

1.

a) AS=(a+Io+Go+Xo-Imo+b(Tro-Txo))=800+3000+4000+1700-200+0.165Q+0.95(7000-1000+0.3Q)=15000+0.65q+0.285Q=15000+0.45Q

b)"SM=\frac{1}{1-((1-0.3)\times0.95-0.165))}=\frac{1}{0.5}=2"

"Q=AS\times SM=(15000+0.45Q)2=30 000+0.9Q"

1-0.9Q=30 000

Q=300 000

c)Qd=Q-Tx+Tr=300 000-(1000+0.3Q)+7000=300 000-1000+0.3*300 000+7000=396 000

d)C=a+bQd=800+0.95Qd=800+0.95*300 000=285 800

e)S=Qd-C=300 000-285 800=14 200

f) Im = 200 + 0.165Q=200+0.165*300 000=49 700

g) NX=X-Im=1700-(200 + 0.165Q)=1700-200-0.165Q=1700-200-0.165*300 000=-48 000

h) GE=G+Tr=4000+7000=11 000

e)BD=GE-Tx=11 000-(1000 + 0.3Q)=11 000-1000-0.3*300 000=-80 000

2. 

a) AS=(a+Io+Go+Xo-Imo+b(Tro-Txo))=800+3000+5500+1700-200+0.165Q+0.95(7000-1000+0.3Q)=16 500+0.65q+0.285Q=16 500+0.45Q

b)"SM=\frac{1}{1-((1-0.3)\times0.95-0.165))}=\frac{1}{0.5}=2"

"Q=AS\times SM=(16 500+0.45Q)2=31500+0.9Q"

1-0.9Q=31 500

Q=315 000

c)Qd=Q-Tx+Tr=315 000-(1000+0.3Q)+7000=315 000-1000+0.3*315 000+7000=415 500

d)C=a+bQd=800+0.95Qd=800+0.95*315 000=300 050

e)S=Qd-C=315 000-300 050=14 980

f) Im = 200 + 0.165Q=200+0.165*315 000=52 175

g) NX=X-Im=1700-(200 + 0.165Q)=1700-200-0.165Q=1700-200-0.165*315 000=-50 475

h) GE=G+Tr=5500+7000=12 500

e)BD=GE-Tx=12 500-(1000 + 0.3Q)=12 500-1000-0.3*315 000=-83 000

3.

a) AS=(a+Io+Go+Xo-Imo+b(Tro-Txo))=800+3500+5500+1700-200+0.165Q+0.95(7000-1000+0.3Q)=17 000+0.65q+0.285Q=17 000+0.45Q

b)"SM=\frac{1}{1-((1-0.3)\times0.95-0.165))}=\frac{1}{0.5}=2"

"Q=AS\times SM=(17 000+0.45Q)2=34 000+0.9Q"

1-0.9Q=34 000

Q=340 000

c)Qd=Q-Tx+Tr=340 000-(1000+0.3Q)+7000=340 000-1000+0.3*340 000+7000=448 000

d)C=a+bQd=800+0.95Qd=800+0.95*340 000=323 800

e)S=Qd-C=340 000-323 800=16 200

f) Im = 200 + 0.165Q=200+0.165*340 000=56 300

g) NX=X-Im=1700-(200 + 0.165Q)=1700-200-0.165Q=1700-200-0.165*340 000=-54 600

h) GE=G+Tr=5500+7000=12 500

e)BD=GE-Tx=12 500-(1000 + 0.3Q)=12 500-1000-0.3*340 000=-90 500

4.

a) AS=(a+Io+Go+Xo-Imo+b(Tro-Txo))=800+3500+5500+2200-200+0.165Q+0.95(7000-1000+0.3Q)=17 500+0.65q+0.285Q=17 500+0.45Q

b)"SM=\frac{1}{1-((1-0.3)\times0.95-0.165))}=\frac{1}{0.5}=2"

"Q=AS\times SM=(17 500+0.45Q)2=35 000+0.9Q"

1-0.9Q=35 000

Q=350 000

c)Qd=Q-Tx+Tr=350 000-(1000+0.3Q)+7000=350 000-1000+0.3*350 000+7000=461 000

d)C=a+bQd=800+0.95Qd=800+0.95*350 000=333 300

e)S=Qd-C=350 000-333 300=16 700

f) Im = 200 + 0.165Q=200+0.165*350 000=57 50

g) NX=X-Im=2200-(200 + 0.165Q)=2200-200-0.165Q=1700-200-0.165*350 000=-55 750

h) GE=G+Tr=5500+7000=12 500

e)BD=GE-Tx=12 500-(1000 + 0.3Q)=12 500-1000-0.3*350 000=-93 500